# Christoph Bandt

### v1

@@ 9 Realizing automata as neighbor graphs of IFS
Example 4 (The dog carpet with five-state automaton)
The self-similar set $X$ in Figure 14 is generated by an IFS with five maps and data from the algebraic number field $\mathbb{Q}(\sqrt{-15}).$ We call it dog carpet because the holes look like a pet. In [7, Figure 10] we found a symmetric version of this fractal with five IFS maps and 12 neighbor maps. This $X$ has an incomplete neighbor automaton with only five states. The curious point is that the maps involve rotations around irrational angles so that the pieces have an infinite number of orientations when we extend the construction over the plane. Examples of this type have not been considered in the literature, and no tiling of this type is expected to exist.
@@
Proof. Let $g(z)=\lambda z$ and $f_{k}=g^{-1}h_{k}$ with orientation-preserving isometries $h_{k}$ for $k=1,...,5.$ We determine the data of this IFS from the automaton. We choose the origin and unit point so that
@@
Finally we consider the path from state $p=h_{1}$ to the inverse state $p-=h_{1}^{-1}$ via $q.$ The corresponding equation is $f_{2}^{-1}f_{4}^{-1}h_{1}f_{5}^{2}=h_{1}^{-1}=h_{4}.$ Hence
@@
$\lambda^{2}-3\lambda+6=0\ \mbox{ with solution }\lambda=\frac{1}{2}(3+i\sqrt{15})\ ,\ a=\frac{1}{4}(1+i\sqrt{15})\ .$
Note that $a$ describes the rotation between pieces $X_{3}$ and $X_{1}$ in Figure 14. Since $2a$ is a quadratic integer, $a$ is not a root of unity, and the rotation angle is irrational. $\Box$
@@
Conclusion
This paper is by no means complete. It is a starting point for further investigation. Open problems were listed in Sections 8 and 9. In Section 4 we pointed out that topology-generating automata should be considered in the broader context of finite type symbolic spaces and graph self-similarity. Our assumption of finite equivalence classes should be cancelled and the decision on the size of equivalence classes included into the algorithm of Section 6. Most importantly, our two algorithms for automata of multiple addresses and finite approximation spaces must be programmed efficiently so that they can be applied to larger automata $G.$ Then new spaces can be generated from automata, and complicated self-similar sets as in [6, 8] can be properly analyzed. Today’s computers provide a chance to realize the vision of Mandelbrot [34] and Barnsley [11] expressed 50 years ago: to model natural geometric phenomena like dust, soil, foam, snow, smoke, clouds. Perhaps automata can also help classify tilings and fractal spaces. On the theoretical side, the isomorphism problems indicated at the end of Section 4 seem most important.

### v2

@@ 9 Realizing automata as neighbor graphs of IFS
Example 9.5 (The dog carpet with five-state automaton).
The self-similar set $X$ in Figure 14 is generated by an IFS with five maps and data from the algebraic number field $\mathbb{Q}(\sqrt{-15})$. We call it dog carpet because the holes look like a pet. In [9, Figure 10] we found a symmetric version of this fractal with five IFS maps and 12 neighbor maps. This $X$ has an incomplete neighbor automaton with only five states. The curious point is that the maps involve rotations around irrational angles so that the pieces have an infinite number of orientations when we extend the construction over the plane. Examples of this type have not been considered in the literature, and no tiling of this type is expected to exist.
@@
Proof 9.7.
Let $g(z)=\lambda z$ and $f_{k}=g^{-1}h_{k}$ with orientation-preserving isometries $h_{k}$ for $k=1,\dots,5$. We determine the data of this IFS from the automaton. We choose the origin and unit point so that
@@
Finally, we consider the path from state $p=h_{1}$ to the inverse state $p-=h_{1}^{-1}$ via $q$. The corresponding equation is $f_{2}^{-1}f_{4}^{-1}h_{1}f_{5}^{2}=h_{1}^{-1}=h_{4}$. Hence
@@
$\lambda^{2}-3\lambda+6=0\ \mbox{ with solution }\lambda=\frac{1}{2}(3+i\sqrt{15})\,\ a=\frac{1}{4}(1+i\sqrt{15}).$
Note that $a$ describes the rotation between pieces $X_{3}$ and $X_{1}$ in Figure 14. Since $2a$ is a quadratic integer, $a$ is not a root of unity, and the rotation angle is irrational.
@@
10 Conclusion
This paper is by no means complete. It is a starting point for further investigation. Open problems were listed in Sections 8 and 9. In Section 4 we pointed out that topology-generating automata should be considered in the broader context of finite type symbolic spaces and graph self-similarity. Our assumption of finite equivalence classes should be canceled and the decision on the size of equivalence classes included into the algorithm of Section 6. Most importantly, our two algorithms for automata of multiple addresses and finite approximation spaces must be programmed efficiently so that they can be applied to larger automata $G$. Then new spaces can be generated from automata, and complicated self-similar sets as in [8, 6] can be properly analyzed. Today’s computers provide a chance to realize the vision of Mandelbrot [35] and Barnsley [12] expressed 50 years ago: to model natural geometric phenomena like dust, soil, foam, snow, smoke, clouds. Perhaps automata can also help classify tilings and fractal spaces. On the theoretical side, the isomorphism problems indicated at the end of Section 4 seem most important.
Acknowledgments
This paper is dedicated to Christiane Frougny. Without her, it would not have been written. Many years ago, she invited me to Paris and introduced me to her numeration community which profoundly influenced my direction of research. I am still very grateful to Christiane.